# Bayesian KalmanNet: Quantifying Uncertainty in Deep Learning Augmented Kalman Filter

### v2

@@ I Introduction
Tracking the latent state of a dynamic system is one of the most common task in signal processing [2]. Often, one is required not only to track the state, but also to provide a reliable assessment of the uncertainty in its estimate. For instance, navigation systems must not only provide the instantaneous position but also characterize its accuracy [3]. Uncertainty quantification is inherently provided by classic algorithms, such as the kf and the ekf (ekf), that rely on the ability to faithfully describe the dynamics as a closed-form ss (ss) model. Such model-based algorithms track not only the state, but also compute the error covariance [4].
@@
The discriminative strategy is task-oriented, learning to directly output the state estimate, following the conventional usage of dnn for classification and regression [6]. Existing designs vary in the knowledge they require on the underlying model. For settings with relatively high level of domain knowledge, e.g., when one can approximately characterize the dynamics as a linear Gaussian ss model, dnn can be enhance classic kf by providing correction terms [20] or pre-processing [21] trained alongside the tracking algorithm. In scenarios with only partially known ss models, dnn augmented kf were proposed, converting the kf into a machine learning algorithm [22, 23, 24, 25, 26, 27, 28, 29]. Doing so bypasses the need to estimate the missing aspects of the ss model, learning to track via end-to-end training. In particular, the KalmanNet architecture [22] and its variants [23, 24, 26, 25] substitute the kg (kg) computation of the ekf with a dnn. The resulting algorithm was shown to enable accurate tracking in partially known and complex dynamics, while facilitating combination with pre-processing [30] and downstream state post-processing [31, 32, 33]. However, as such methods learn to estimate the state, they do not naturally provide the error covariance, for which there is rarely "ground-truth".
@@
Our analysis identifies that reliable extraction of the error covariance from existing dnn-aided algorithms is feasible only in some settings, e.g., when tracking fewer state variables than observed ones. Moreover, doing so requires additional knowledge of the ss model that is not required for state tracking. To overcome this, we propose a novel dnn-aided tracking algorithm that supports uncertainty quantification termed Bayesian KalmanNet. Our proposed algorithm draws inspiration from Bayesian deep learning [37, 38, 39], deviating from conventional frequentist learning to augment an ekf with a stochastic dnn to compute the kg. Doing so enables tracking the with the same level of partial ss modeling needed by conventional KalmanNet and its variants, while providing reliable uncertainty measures by sampling multiple dnn realizations.
@@ II-E Bayesian Deep Learning
Bayesian deep learning offers a framework to quantify uncertainty in dnn predictions. It treats the dnn parameters as random variables, and learns their posterior distribution given the data. This probabilistic approach provides a measure of uncertainty by sampling multiple realizations of the (random) dnn during inference [45].
@@
To learn a stochastic dnn, one typically seeks a distribution on the weights, i.e., $p({\boldsymbol{\theta}}|\mathcal{D})$, that is dictated by some parameters ${\boldsymbol{\phi}}$. This can result in either an explicit formulation of the distribution, as in variational inference [46], or via alternative induced trainable stochasticity, as in Monte Carlo dropout [47]. Once dictated, the parameters of the dnn distribution, ${\boldsymbol{\phi}}$, are trained using the elbo (elbo) objective, which regularizes the expected loss with a term minimizing the kl (kl) divergence between the induced distribution $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$ and some pre-determined prior $p({\boldsymbol{\theta}})$, i.e.,
$\text{ELBO}_{\mathcal{D}}({\boldsymbol{\phi}})=\mathbb{E}_{q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})}\left[\log p(\mathcal{D}|{\boldsymbol{\theta}})\right]-\text{KL}(q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})|p({\boldsymbol{\theta}})).$ (12)
@@ III-B1 Empirical Error Variance
the loss includes an additive term that compares the the predicted error with its empirical valuate obtained from (18). By defining the second moment loss over the whole dataset as:
@@
where $\beta\in[0,1]$ balances the contribution of each term. The hyperparameters $\beta$ can be gradually increased as training progresses to first learn to track and then also encourage reliable error prediction. Specifically, when setting $\beta=0$, then training with (20) coincides with training with the standard $\ell_{2}$ loss.
@@ III-B2 Gaussian Prior
We note that when the error covariance $\hat{\boldsymbol{\Sigma}}_{t}$ is not estimated, but is instead assumed to be a scaled identity, then training with the loss in (III-B2) is equivalent to training with the standard $\ell_{2}$ loss.
@@ III-C Discussion
We identified three different approaches to extract error covariance in discriminative dnn-aided tracking in non-linear ss models. These approaches vary based on the architecture employed and the level of interpretability it provides. It is emphasized though that for all considered approaches, one must still alter the learning procedure for the extracted features to be considered as uncertainty measures. To encourage learning to accurately predict the error, we propose two training objectives. This examination of error covariance extraction in dnn-aided kf combines new approaches, while encompassing techniques used for specific schemes (e.g., (III-B2) used by rkn [9]), and extending previous findings (e.g., (17) extends [36] to non-linear models).
@@ IV-A Bayesian KalmanNet Tracking Algorithm
Uncertainty-Aware Tracking: The proposed augmentation of the ekf with Bayesian dnn allows to track as in KalmanNet, while leveraging ensembling to predict the estimation error. When tracking is launched, we sample $J$ i.i.d. realizations of the kg computations dnn from $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$, denoted $\{{\boldsymbol{\theta}}_{j}\}_{j=1}^{J}$, with $J$ being a hyperparameter. For instance, when employing Monte Carlo dropout, each dnn realization is obtained from the initial fixed weight ${\boldsymbol{\theta}}_{j}$ and a realization of the dropout masks sampled with multivariate Bernoulli distribution $\boldsymbol{p}$.
@@
The resulting operation is summarized as Algorithm 1 (in which, for brevity, we write the output of the $j$th kg dnn realization at time $t$ as $\hat{\mathbf{K}}^{j}_{t}$), and illustrated in Fig. 2.
@@ IV-B Training Bayesian KalmanNet
We follow the specialization of the elbo objective to concrete Monte Carlo based dnn proposed in [47], while altering the data matching term to incorporate explicit encouraging of reliable error variance recovery, as proposed in Subsection III-B for frequentist learning. To formulate this, let $L$ denote the number of layers in the dnn in which dropout is incorporated, and let $\{\lambda_{l}\}_{l=1}^{L}$ denote their indices. In the implementation detailed in Subsection IV-A, these indices correspond to the fc layer of KalmanNet’s Architecture 1. We write the parameters of layer $\lambda$ of ${\boldsymbol{\theta}}$ as ${\boldsymbol{\theta}}^{\lambda}$ and its dropout coefficient as $p^{(\lambda)}$. The proposed loss function, based on the labeled data in (8), is formulated as
@@
where $\mathcal{L}_{\lambda_{l}}^{\rm DO}({\boldsymbol{\phi}})$ computes the kl-based regularization term of the layer $\lambda_{l}$ based on the learned multivariate Bernoulli distribution, with pre-determined prior in (12) being a maximal entropy Bernoulli distribution. The layer-wise loss term is given by [47]
@@
In (24), $n^{(\lambda)}$ is the number of input features to the $\lambda$th layer; $c_{1}$ and $c_{2}$ are constant coefficients (hyperparameters); and $\mathcal{H}(p)$ is the entropy of a Bernoulli random variable with probability $p$, defined as
@@
This loss in (23) guides the tuning of ${\boldsymbol{\phi}}$ when training Bayesian KalmanNet to both extract an accurate state estimation and error covariance (by taking the expectation of our proposed empirical error variance as the data matching term), while encouraging randomness in the prior on the dropout probability $\boldsymbol{p}$.
TABLE I: Domain knowledge requirements comparison
@@ IV-C Discussion
The incorporation of Bayesian deep learning techniques can be viewed as trading computational complexity for uncertainty measures. This induced complexity also translates into additional inference latency, as reported in Section V. Specifically, the application of Bayesian KalmanNet can be viewed as $J$ realizations of KalmanNet with different parameterization, thus increasing the computational burden by a factor of $J$ at each time instance. In that sense, Bayesian KalmanNet bears similarity to alternative tracking algorithm that implement multiple Kalman-type filters in parallel, e.g., [48], though the conventional approach uses a different ss model with deterministic parameters for each filter, while Bayesian KalmanNet uses multiple stochastic realizations of the filter obtained for the same (partially known) ss model, following the task-oriented discriminative learning paradigm [13].
@@
Our design of Bayesian KalmanNet gives rise to numerous possible avenues for extension. While our formulation accommodates a generic Bayesian dnn, our implementation and training objective focus on the usage of Monte Carlo dropout. Our formulation differs than standard Bayesian dnn as all realizations are based on the same prior. One can possibly consider alternative Bayesian parameterizations, based on, e.g., variational inference with explicit priors [46], as well as its incorporation with different KalmanNet-type architectures, e.g., [23, 24, 26, 25]. Moreover, the combination of Bayesian deep learning and KalmanNet indicates that such methodologies can also be employed in alternative forms of model-based deep learning [12]. We leave these extensions for future work.
@@ V-A1 Tracking Algorithms
Unless stated otherwise, we use (15) to convert the kg and prior covariance into error covariance. We train rkn and skn with (III-B2); KalmanNet is trained with (9); while Bayesian KalmanNet is trained using (23). All data-driven algorithms are trained with the same data comprised of $|\mathcal{D}|=300$ trajectories.
@@ V-A2 Data generation
Unless stated otherwise, throughout the follow experimental study, we set the observation noise covariance $\mathbf{R}$ to take the form $r^{2}\mathbf{I}$, with $1/r^{2}=n/{\rm Tr}(\mathbf{R})$ denoting the snr (snr). The ratio between the state evolution and the observation noise variances is fixed to $0.01$, such that changes in the snr affect both the state evolution and the observation noises. To perform some statistical tests and evaluation, mc simulation was performed, with $N=100$ trajectories per scenario.

### v3

@@ I Introduction
Tracking the latent state of a dynamic system is one of the most common tasks in signal processing [2]. Often, one is required not only to track the state, but also to provide a reliable assessment of the uncertainty in its estimate. For instance, navigation systems must not only provide the instantaneous position but also characterize its accuracy [3]. Moreover, uncertainty quantification in state estimation provides a measure of confidence in the predictions and enables informed, risk-aware decision-making, which is particularly critical in safety-sensitive domains. Uncertainty quantification is inherently provided by classic algorithms, such as the kf and the ekf (ekf), that rely on the ability to faithfully describe the dynamics as a closed-form ss (ss) model. Such model-based algorithms track not only the state, but also compute the error covariance [4].
@@
The discriminative strategy is task-oriented, learning to directly output the state estimate, following the conventional usage of dnn for classification and regression [6]. Existing designs vary in the knowledge they require on the underlying model. For settings with relatively high level of domain knowledge, e.g., when one can approximately characterize the dynamics as a linear Gaussian ss model, dnn can enhance classic kf by providing correction terms [20] or pre-processing [21] trained alongside the tracking algorithm. In scenarios with only partially known ss models, dnn augmented kf were proposed, converting the kf into a machine learning algorithm [22, 23, 24, 25, 26, 27, 28, 29]. Doing so bypasses the need to estimate the missing aspects of the ss model, learning to track via end-to-end training. In particular, the KalmanNet architecture [22] and its variants [23, 24, 26, 25] substitute the kg (kg) computation of the ekf with a dnn. The resulting algorithm was shown to enable accurate tracking in partially known and complex dynamics, while facilitating combination with pre-processing [30] and downstream state post-processing [31, 32, 33]. However, as such methods learn to estimate the state, they do not naturally provide the error covariance, for which there is rarely "ground-truth".
@@
Our analysis identifies that reliable extraction of the error covariance from existing dnn-aided algorithms is feasible only in some settings, e.g., when tracking fewer state variables than observed ones. Moreover, doing so requires additional knowledge of the ss model that is not required for state tracking. To overcome this, we propose a novel dnn-aided tracking algorithm that supports uncertainty quantification termed Bayesian KalmanNet. Our proposed algorithm draws inspiration from Bayesian deep learning [37, 38, 39], deviating from conventional frequentist learning to augment an ekf with a stochastic dnn to compute the kg. Doing so enables tracking the state with the same level of partial ss modeling needed by conventional KalmanNet and its variants, while providing reliable uncertainty measures by sampling multiple dnn realizations.
@@ II-E Bayesian Deep Learning
While the formulation of the ss model in Subsection II-A and the model-based tracking in Subsection II-B focuses on a Bayesian setting, from a deep learning perspective, the term Bayesian deep learning refers to a framework in which a stochastic modeling is employed on the dnn itself. Accordingly, we henceforth use the term frequentist within the context of dnn to refer to the standard deep learning approach in which the weights of a dnn are deterministic weights, whereas a Bayesian dnn models its weights as random variables governed by a learned posterior distribution [37].
Specifically, Bayesian deep learning offers a framework to quantify uncertainty in dnn predictions. It treats the dnn parameters as random variables, and learns their posterior distribution given the data. This probabilistic approach provides a measure of uncertainty by sampling multiple realizations of the (random) dnn during inference [45].
@@
To learn a stochastic dnn, one typically seeks to approximate a posterior distribution on the weights, i.e., $p({\boldsymbol{\theta}}|\mathcal{D})$, that is dictated by some parameters ${\boldsymbol{\phi}}$. This can result in either an explicit formulation of the distribution, as in variational inference [46], or via alternative induced trainable stochasticity, as in Monte Carlo dropout [47]. Once dictated, the parameters of the dnn distribution, ${\boldsymbol{\phi}}$, are trained using the elbo (elbo) objective, which regularizes the expected loss with a term minimizing the kl (kl) divergence between the induced distribution $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$ and some pre-determined prior $p({\boldsymbol{\theta}})$, i.e.,
$\text{ELBO}_{\mathcal{D}}({\boldsymbol{\phi}})=\mathbb{E}_{q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})}\left[\log p({\boldsymbol{\theta}}|\mathcal{D})\right]-\text{KL}(q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})|p({\boldsymbol{\theta}})).$ (12)
@@ III-B1 Empirical Error Variance
the loss includes an additive term that compares the predicted error with its empirical value obtained from (18). By defining the second moment loss over the whole dataset as:
@@
The hyperparameter $\beta\in[0,1]$ in (20) is used to balance the trade-off between minimizing the tracking error and encouraging accurate prediction of the error covariance. In our numerical implementation detailed in Section V, $\beta$ is initialized with a small value at the beginning of training, so that the network initially focuses on learning to accurately track the state. As training progresses, $\beta$ is gradually increased, thereby placing more emphasis on the accuracy of the predicted uncertainty. Specifically, when setting $\beta=0$, then training with (20) coincides with training with the standard $\ell_{2}$ loss.
@@ III-B2 Gaussian Prior
We note that when the error covariance $\hat{\boldsymbol{\Sigma}}_{t}$ is not estimated, but is instead assumed to be a scaled identity, then training with the loss in (21) is equivalent to training with the standard $\ell_{2}$ loss.
@@ III-C Discussion
We identified three different approaches to extract error covariance in discriminative dnn-aided tracking in non-linear ss models. These approaches vary based on the architecture employed and the level of interpretability it provides. It is emphasized though that for all considered approaches, one must still alter the learning procedure for the extracted features to be considered as uncertainty measures. To encourage learning to accurately predict the error, we propose two training objectives. This examination of error covariance extraction in dnn-aided kf combines new approaches, while encompassing techniques used for specific schemes (e.g., (21) used by rkn [9]), and extending previous findings (e.g., (17) extends [36] to non-linear models).
@@ IV-A Bayesian KalmanNet Tracking Algorithm
Uncertainty-Aware Tracking: The proposed augmentation of the ekf with Bayesian dnn allows to track as in KalmanNet, while leveraging ensembling to predict the estimation error. When tracking is launched, we sample $J$ i.i.d. realizations of the kg computations dnn from $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$, denoted $\{{\boldsymbol{\theta}}_{j}\}_{j=1}^{J}$, with $J$ being a hyperparameter. The type of distribution represented by $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$ is a design hyperparameter, as different families of distributions are considered in the Bayesian deep learning literature [46]. In our numerical study reported in Section V we employ Monte Carlo dropout [47], where each dnn realization is obtained from the initial fixed weight ${\boldsymbol{\theta}}_{j}$ and a realization of the dropout masks sampled with multivariate Bernoulli distribution $\boldsymbol{p}$. However, the formulation of Bayesian KalmanNet is not restricted to Monte Carlo dropout, and can be combined with alternative distributions, e.g., Gaussian weights whose parameters are the first- and second-order moments.
@@
From a Bayesian deep learning perspective, the matrix $\hat{\boldsymbol{\Sigma}}_{t}$ in (22b) represents an empirical estimate of the predictive uncertainty, namely, the second-order moment of the posterior predictive distribution over the state $\mathbf{x}_{t}$ given the observations and training data [48]. Specifically, each estimate $\hat{\mathbf{x}}_{t}^{(j)}$ in (22a) corresponds to a sample from a different model instantiation, i.e., from the distribution $q({\boldsymbol{\theta}}|{\boldsymbol{\phi}})$ over the neural network weights. This ensemble of predictions approximates the posterior predictive distribution, and the sample covariance $\hat{\boldsymbol{\Sigma}}_{t}$ estimates the total predictive uncertainty. While there is no general guarantee that individual estimators $\hat{\mathbf{x}}_{t}^{(j)}$ are realizations of an unbiased estimator, the usage of its sample covariance for quantifying uncertainty is a standard and widely accepted approach in Bayesian deep learning. Accordingly, our proposed Bayesian KalmanNet uses $\hat{\boldsymbol{\Sigma}}_{t}$ in (22b) as an estimate of the error covariance. The resulting operation is summarized as Algorithm 1 (in which, for brevity, we write the output of the $j$th kg dnn realization at time $t$ as $\hat{\mathbf{K}}^{j}_{t}$), and illustrated in Fig. 2.
@@ IV-B Training Bayesian KalmanNet
We follow the specialization of the elbo objective to concrete Monte Carlo based dnn proposed in [47], while altering the data matching term to incorporate explicit encouraging of reliable error variance recovery, as proposed in Subsection III-B for frequentist learning. To formulate this, let $L$ denote the number of layers in the dnn in which dropout is incorporated, and let $\{\lambda_{l}\}_{l=1}^{L}$ denote their indices. In the implementation detailed in Subsection IV-A, these indices correspond to the fc layer of KalmanNet’s Architecture 1. We write the parameters of layer $\lambda$ of ${\boldsymbol{\theta}}$ as ${\boldsymbol{\theta}}^{\lambda}$ and its dropout coefficient as $p^{(\lambda)}$. Accordingly, the trainable parameters of Bayesian KalmanNet in such settings, denoted ${\boldsymbol{\phi}}$, include the weights of each of the layers ${\boldsymbol{\theta}}_{0}$, as well as the distribution parameters for the layers in which dropout is incorporated, denoted by $\boldsymbol{p}=[p^{\lambda_{1}},\ldots,p^{\lambda_{L}}]$. The proposed loss function, based on the labeled data in (8), is formulated as
@@
where $\mathcal{L}_{\lambda_{l}}^{\rm KL}({\boldsymbol{\phi}})$ computes the kl-based regularization term of the layer $\lambda_{l}$ based on the learned multivariate Bernoulli distribution, with pre-determined prior in (12) being a maximal entropy Bernoulli distribution. The empirical loss term and the KL loss term in (23) are balanced through the non-negative hyperparameters $c_{1},c_{2}$, that appear in the layer-wise loss term, which is given by [47]
@@
In (24), $n^{(\lambda)}$ is the number of input features to the $\lambda$th layer; and $\mathcal{H}(p)$ is the entropy of a Bernoulli random variable with probability $p$, defined as
@@
This loss in (23) guides the tuning of ${\boldsymbol{\phi}}$ when training Bayesian KalmanNet to both extract an accurate state estimation and error covariance (by taking the expectation of our proposed empirical error variance as the data matching term), while encouraging randomness in the prior on the dropout probability $\boldsymbol{p}$. Having formulated the loss in (23), the training of Bayesian KalmanNet follows the conventional stochastic gradient based optimization of dnn, only with a Bayesian loss term. Accordingly, the overall training complexity is comparable to that of frequentist KalmanNet [22].
TABLE I: Domain knowledge requirements comparison.
@@ IV-C Discussion
The uncertainty measure extracted by Bayesian KalmanNet encapsulates the overall estimation error covariance. This uncertainty measure inherently comprises two components [48]: aleatoric uncertainty, resulting from the stochasticity of the system (e.g., process and measurement noises); and epistemic uncertainty, arising due to limited or mismatched knowledge within the estimator, particularly when the underlying dynamics are only partially known or when trained with a limited dataset. Classical model-based methods, such as the ekf, can reliably quantify aleatoric uncertainty when operating with an accurate ss model and well-characterized noise statistics, they underperform when the model is misspecified. Conversely, existing frequentist dnn-aided kf struggle to provide any form of uncertainty quantification, even with abundant training data. The approaches proposed for extending frequentist dnn-aided kf introduced in Section III are geared towards quantifying aleatoric uncertainty, aiming to recreate the covariance computations of the model-based ekf without explicitly propagating second-order moments. Accordingly, the resulting characterization, which does not encompass epistemic uncertainty, is expected to be sensitive to the amount of training data. Bayesian KalmanNet overcomes these limitations by providing uncertainty estimates that reflect both sources of uncertainty, thereby yielding more reliable and calibrated error covariances under data scarcity and model mismatch. This dual capability makes Bayesian KalmanNet particularly suitable for practical scenarios where both the dynamics and noise statistics are only partially known.
@@
The incorporation of Bayesian deep learning techniques can be viewed as trading computational complexity for uncertainty measures. This induced complexity also translates into additional inference latency, as reported in Section V. The hyperparameter $J$ plays a key role in affecting both the performance, as increasing $J$ enhances the approximation of the posterior predictive distribution, while also increasing complexity. Specifically the application of Bayesian KalmanNet can be viewed as $J$ realizations of KalmanNet with different parameterizations, thus increasing the computational burden by a factor of $J$ at each time instance. The induced latency of Bayesian KalmanNet can be notably reduced using parallel computing, as each of the $J$ realizations can be applied simultaneously. We also note that this operation of Bayesian KalmanNet bears similarity to alternative tracking algorithm that implement multiple Kalman-type filters in parallel, e.g., [49], though the conventional approach uses a different ss model with deterministic parameters for each filter, while Bayesian KalmanNet uses multiple stochastic realizations of the filter obtained for the same (partially known) ss model, following the task-oriented discriminative learning paradigm [13].
@@
Our design of Bayesian KalmanNet gives rise to numerous possible avenues for extension. Our design, as well as that considered for frequentist KalmanNet [22], assumes that the ss model in (1) does not change. Thus, our approach is not naturally transferable to time-varying dynamics, for which adaptive kf, often based on variational inference [50, 51, 52], are required. Previous works that considered the KalmanNet methodology in time-varying dynamics, which proposed adaptation based on hypernetworks [53] and unsupervised learning [54], focused on block-wise variations in the ss model. The extension of the KalmanNet methodology to time-varying dynamics is the area of ongoing research, which, inspired by recent advances in continual Bayesian learning [55, 56], can potentially benefit from the incorporation of Bayesian deep learning. Moreover, while our formulation accommodates a generic Bayesian dnn, our implementation detailed in Section V is based on an architecture proposed in [22], and training objective focus on the usage of Monte Carlo dropout. Our formulation differs than standard Bayesian dnn as all realizations are based on the same prior. One can possibly consider alternative Bayesian parameterizations, based on, e.g., variational inference with explicit priors [46], as well as its incorporation with different KalmanNet-type architectures, e.g., [23, 24, 26, 25]. Moreover, the combination of Bayesian deep learning and KalmanNet indicates that such methodologies can also be employed in alternative forms of model-based deep learning [12]. We leave these extensions for future work.
@@ V-A1 Tracking Algorithms
Unless stated otherwise, we use (15) to convert the kg and prior covariance into error covariance. We train rkn and skn with (21); KalmanNet is trained with (9); while Bayesian KalmanNet is trained using (23). All data-driven algorithms are trained with the same data comprising of $|\mathcal{D}|=300$ trajectories. Training commences with the dnn weights initialized via standard Xavier initialization, and the Monte Carlo dropout parameters of BKN randomized uniformly in $[0.5,0.8]$, following the common practice in Bayesian dnn [47]. The remaining learning hyperparameters were selected based on standard cross validation.
@@ V-A2 Data generation
Unless stated otherwise, throughout the following experimental study, we set the observation noise covariance $\mathbf{R}$ to take the form $r^{2}\mathbf{I}$, with $1/r^{2}=n/{\rm Tr}(\mathbf{R})$ denoting the snr (snr). The ratio between the state evolution and the observation noise variances is fixed to $0.01$, such that changes in the snr affect both the state evolution and the observation noises. All dnn-aided algorithms were trained on a range of snr rather than for each snr separately, to emphasize their robustness and ability to operate across different noise levels. To perform statistical tests and evaluation, mc simulation was performed, with $N=100$ trajectories per scenario.